OpenStudy (rainbow_dashie):

Simplify the following expression. (43x + 63y) + (36x - 21y) 79x + 84y 106x + 84y 79x + 42y 106x + 42y

OpenStudy (anonymous):

43X+36= 63Y-21Y=

OpenStudy (anonymous):

it's C

OpenStudy (rainbow_dashie):

ive never heard of this

OpenStudy (cicilybailey14):

Let's simplify step-by-step. 43x+63y+36x−21y =43x+63y+36x+−21y Combine Like Terms: =43x+63y+36x+−21y =(43x+36x)+(63y+−21y) =79x+42y Answer: =79x+42y

OpenStudy (cicilybailey14):

http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

43X+36X+63Y-21y=79X+42Y

OpenStudy (anonymous):

A

OpenStudy (anonymous):

IT IS C

OpenStudy (rainbow_dashie):

ay which 1

OpenStudy (anonymous):

Sorry it c

OpenStudy (cicilybailey14):

lmao it's c

OpenStudy (rainbow_dashie):

Simplify the following expression. (19x - 17y) + (47x + 65y) 2x + 48y 2x + 82y 66x + 82y 66x + 48y

OpenStudy (anonymous):

19x+47=66x 17y+65y=82y so 66x+82y C

OpenStudy (rainbow_dashie):

it was d

OpenStudy (rainbow_dashie):

Simplify the following expression. (20x + 7y) + (39x + 34y) 27x + 41y 27x + 54y 59x + 54y 59x + 41y

OpenStudy (rainbow_dashie):

@1497

OpenStudy (anonymous):

20x+39x=59x 7y+34y=41y 59x+41y d

OpenStudy (rainbow_dashie):

Simplify the following expression. (33x - 26y) + (45x + 59y) 7x + 33y 7x + 85y 78x + 33y 78x + 85y @1497

OpenStudy (anonymous):

wow that's a lot of question

OpenStudy (rainbow_dashie):

ik

OpenStudy (anonymous):

33x+45x=78x 26y+59y=85y 78x+85y d

OpenStudy (anonymous):

no it's C:

OpenStudy (anonymous):

78X+33Y

OpenStudy (anonymous):

that's 111

OpenStudy (anonymous):

59Y-26Y=33Y

OpenStudy (anonymous):

THE ANS is C:

OpenStudy (anonymous):

http://www.acalculator.com/math-calculators.html